For the male stature:

- sample population mean = 66.5"

- standard deviation = 1.3"

What male stature for this population represents the :

- 10th percentile?

- 50th percentile?

- 95th percentile?
Answer #1

Solution:

We are given µ = 66.5, σ = 1.3

For 10th percentile, Z = -1.28155 (by using z-table)

10th percentile = µ + Z*σ

10th percentile = 66.5 - 1.28155*1.3

10th percentile = 64.83399

For 50th percentile, Z = 0 (by using z-table)

50th percentile = µ + Z*σ = 66.5 + 0*1.3 = 66.5

50th percentile = 66.5

For 95th percentile, Z = 1.644853627 (by using z-table)

95th percentile = µ + Z*σ = 66.5 + 1.644853627*1.3 = 68.63830972

95th percentile = 68.63830972